# Ego-Exo4D: Understanding Skilled Human Activity from First- and Third-Person Perspectives

### v3

@@ 13.A.2 Ego-Exo translation
In Table 12 we provide a break down of the results across different scenarios, using GNT-mask for track prediction and DiT-pix for clip generation. We can observe similar trends for the two subtasks: the methods achieve better results in basketball and soccer scenarios than in bike and cook scenarios, which is reasonable as the objects in bike and cook scenarios are more complex and diverse.
Contributions statement Lorenzo Torresani co-led the translation benchmark, developed the task formulation, advised the baseline development, and contributed to editing this section. Judy Hoffman co-led the translation benchmark and advised the baseline development. Feng Cheng led the baseline development and implemented the pix2pix and DiT models for track prediction and clip generation. Mi Luo implemented the GNT baseline and the evaluation pipeline for ego track prediction. Ziwei Zhao contributed the pix2pix baseline for multi-frame input and led the evaluation for ego clip generation. Huiyu Wang advised the baseline development and contributed to the task definition, the baseline design, and the metric selection and analysis.
Figure 31: Qualitative ego clip generation by DiT-pix on the test set. The model takes 6 input images (exo frame, exo crop, exo mask, exo crop mask, ego mask, and ego crop mask). Note that only the exo frame and the exo crop images are included in this visualization. The ego frame, serving solely as a reference, does not constitute either an input or an output element.
Figure 32: Qualitative demonstration of the importance of the different inputs given to DiT-pix. The exo crop image and ego crop mask are critical for good performance.
Figure 33: Comparison of ego clip generations using frame-to-frame vs clip-to-clip variants of DiT-pix. The clip-to-clip version of the model produces outputs that are more coherent across the frames of the clip, even for frames where the exo view is heavily occluded.
@@
 Basketball 16.2/10.4 0.38/0.36
Soccer 17.0/6.6 0.45/0.51
Music 4.3/9.0 0.28/0.18
Health 11.7/7.8 0.41/0.21
Bike 5.9/5.7 0.51/0.43
Cook 9.5/6.0 0.49/0.41
Table 12: Breakdown of results per scenario for the subtasks of ego track prediction (IoU) and ego clip generation (LPIPS) on the v1/v2 test set, respectively. All models are trained on the v1 training set.
13.B   Ego-(Exo) Keystep Recognition
This family of tasks centers around recognizing the keysteps of a procedural activity and modeling their dependencies. Specifically, there are three tasks: fine-grained keystep recognition (Sec. 13.B.1), efficient multimodal keystep recognition (Sec. 13.B.2), and procedure understanding (Sec. 13.B.3). We refer to the family of tasks as “ego-(exo)" since exo may be available at the time of training but not inference.
13.B.1 Fine-grained keystep recognition
Annotations
@@ 13.B.1 Fine-grained keystep recognition
We annotate videos featuring any of the three procedural activities (i.e., cooking, bike repair, health) with temporal segments of keysteps, i.e., actions that contribute towards the completion of a procedural task. To accurately model the hierarchical nature of the activities, we also develop a data-driven hierarchical keystep taxonomy concurrently with the annotation process.
Figure 34 shows the annotation user interface. We provide annotators a composite view of time-synchronized ego and exo videos. Each keystep annotation contains the start and end timestamps, a category label, a natural language description, and a flag indicating whether the keystep is essential or optional for task completion. Annotators interact with a search widget which displays keystep labels with their complete path within a hierarchical tree, e.g., Making cucumber & tomato salad $>$ Prepare dressing $>$ To a bowl or jar $>$ Add salt.
@@
Formal task definition
Classification of trimmed video clips is a problem formulation commonly adopted in action recognition benchmarks [67, 149, 46]. However, our task differs from action recognition in two fundamental aspects. First, it targets fine-grained keystep recognition rather than classification of coarse activities. We note that this adds significant complexity, since different keysteps of an activity often involve manipulating the same objects in the scene (e.g., folding the bedsheet and smoothing out the bedsheet) and are consequently difficult to tell apart. Furthermore, different keysteps may be represented over largely different time spans (e.g., the average time span for “kneading dough” is 87.3 seconds, in stark contrast with “getting salt”, which averages at 3.6 seconds), thus requiring analysis at different levels of temporal granularity. The second key difference is the potential to leverage contextual cues available in exocentric videos during training to improve the prediction accuracy on egocentric videos. Note that at test time, the input to the model includes just the ego-view videos (RGB only). Exo-view videos, activity and scenario names, narrations, audio and associated metadata such as eye gaze, 3D point clouds, camera pose, and IMU information are excluded as inputs for inference (although we encourage exploring their potential utility in training) as our ultimate goal is a vision-centric approach that performs egocentric keystep recognition.
Baselines
  • Action classification. As a prototypical example of this classic genre, we select a TimeSformer [13] model initialized with the checkpoint pretrained on the large-scale third-person action dataset Kinetics-600 [67] due to its strong performance in various video understanding tasks.
  • Video-language pretraining. We adopt the EgoVLPv2 framework [123] and initialize the model with two backbones, one pretrained on the Ego4D dataset [47] (which contains only ego views) and the other that is further pretrained on EgoExo4D (which encompasses both ego and exo views), to exploit the variances in pretraining between these datasets.
  • View-invariant learning. A two-stage training approach is employed. In the first stage, we utilize all available (ego, exo) video pairs in the dataset for training a view-invariant (VI) encoder. The training objective is a clip-level contrastive loss [110], aiming at identifying the synchronized (ego, exo) pairs as positive, and the non-synchronized pairs as negative. In the second stage, this pretrained model is further trained with a classification loss, aligning with the clip-level classification nature of the downstream task. Note that to align with the clip-level classification task, our contrastive loss operates at the clip-level, rather than at the frame-level as was done in view-invariant loss proposed in [140, 144].
  • Viewpoint distillation. This also adopts a two-stage training approach. In the first stage, we train a multi-view teacher that takes both ego and exo views as input. In the second stage, a single-view ego student is trained, distilling knowledge [52] from the multi-view teacher to encapsulate information from both views.
  • Ego-Exo Transfer. Here we follow the methodology proposed in Ego-Exo [82] which uses egocentric pseudo-labels to pre-train the network. We employ a masked autoencoder (MAE) [162] backbone, initialized from a Kinetics checkpoint, and the pseudo-labels provided from the Ego-Exo checkpoint to fine-tune with two auxiliary heads (Object-Score and Interaction-Map). We then further finetune the model with a classification loss for fine-grained keystep recognition.
For the first two baselines (which utilize pretrained checkpoints from well-established benchmarks), two training settings are further examined: one using only the ego view for classification loss and the other utilizing both ego and exo view videos, with the training objective being the sum of ego view and exo view classification losses.
@@
Data
The keysteps in our dataset exhibit a very long-tailed distribution. To address this challenge, we set a cutoff threshold at 20 samples per keystep, limiting our analysis to 278 unique keysteps as shown in Figure 36. For simplicity, we consider only the leaf node keysteps in the hierarchy. Exploring the hierarchical structure including parent nodes is a promising direction but we leave this as future work. In all, the dataset for keystep recognition comprises 129,91/130,979 segments in the v1/v2 data release, respectively, with an average duration of 11.34 seconds each. Specifically, the training set contains 74,457/74,342 segments, of which 14,550/14,326 are from the ego view and the rest from the exo view for v1/v2, respectively. The validation set consists of 23,088/23,636 segments, including 4,502/4,517 ego-view segments, and the test set has 32,369/33,001 segments with 6,351/6,373 in the ego view for v1/v2, respectively.
Ego Accuracy (%)
Method Train data (v1) Val (v1/v2) Test (v1/v2)
TimeSFormer [13] (K600) ego 35.25/35.13 35.24/35.93
TimeSFormer [13] (K600) ego,exo 32.67/32.68 29.84/31.04
EgoVLPv2 [123] (Ego4D) ego 36.89/36.51 37.51/37.55
EgoVLPv2 [123] (Ego4D) ego,exo 37.03/35.84 36.84/36.59
EgoVLPv2 [123] (EgoExo) ego 37.61/36.04 37.85/37.72
EgoVLPv2 [123] (EgoExo) ego,exo 38.21/39.10 38.69/38.76
VI Encoder [110] (EgoExo) ego,exo 40.23/40.34 40.61/41.53
Viewpoint Distillation [52] ego,exo 37.79/38.19 38.10/39.49
Ego-Exo Transfer MAE [82] ego,exo 36.71/37.17 35.57/36.58
Table 14: Top-1 accuracy of keystep recognition on the v1/v2 data release, respectively. All models are trained on keystep v1 data. The pre-training dataset is denoted in parentheses. VI is short for view-invariant.
Results
Table 14 reports the Top-1 accuracy for ego classification on both validation and test sets. Among all the baselines, the VI Encoder emerges as the top performer, achieving a v2 test accuracy of 41.53%. It is closely followed by Viewpoint Distillation and EgoVLPv2 pretrained on EgoExo, which attain v2 test accuracies of 39.49% and 38.76% respectively. These results open discussion on how to effectively utilize exo videos during training to enhance ego keystep recognition during test time.
First, we note that different approaches respond differently to the addition of exo-view videos during training. Specifically, while the TimeSFormer (K600) exhibits a degradation when the exo classification loss is integrated into the objective (i.e., v2 test accuracy drops from 35.93% to 31.04%, EgoVLPv2 pretrained on EgoExo benefits from the introduction of exo-view videos (i.e., v2 test accuracy improves from 37.72% to 38.76%. This enhancement is also evident in the VI encoder and viewpoint distillation when compared to TimeSFormer (K600) that only utilizes ego-view videos for training. These observations suggest that certain baselines are better equipped at leveraging exo information during training to improve ego keystep recognition.
@@ Results
We follow with a more detailed analysis of per-step performance in Figure 37, comparing training with ego-view videos and exo-view videos. We can observe that exo views show performance advantages over ego views in several steps, with the keystep ‘have a conversation asking different questions’ benefiting the most from exo. Conversely, ego views are more effective in steps involving manipulation of small objects, like ‘cut carrots’ and ‘unpack the new tube’. This observation can be linked to the positioning of exo cameras, which are often placed further away from the subject, enabling them to capture a broader view, though possibly missing finer details of the activity. We hope these findings provide insight for future research on the effective use of exo-view videos during training.
Overall, we posit that the endeavor to enhance view-invariant learning and to more effectively harness the complementary information from exo views for ego keystep recognition remains an open avenue. Our findings underscore the need for further investigation and innovation in this domain.
@@
Figure 37: Keystep recognition evaluated per keystep label on the v1 validation set, comparing training with only ego-view videos versus exo-view videos. The accuracy delta (exo-ego) is displayed, where a positive value indicates better exo and a negative value indicates better ego performance.
Contribution statement
Tushar Nagarajan co-led the keystep recognition benchmark, co-developed the task formulation, and advised the baseline development. David Crandall co-led the keystep recognition benchmark and contributed to the task formulation. Yale Song co-led the keystep recognition benchmark and led the keystep annotation effort, including design of annotation guidelines, taxonomy development and coordination of annotation workflows; he also contributed to the task formulation, advised baseline design, and facilitated the delivery of EgoVLPv2 pretrained backbone. Triantafyllos Afouras contributed to the taxonomy definition, managed the labeling effort, and developed software for post-processing the annotations. Zihui Xue led the baseline development effort, implemented the TimeSFormer, EgoVLP, VI Encoder and Viewpoint distillation baselines, and performed analysis of results. Eugene Byrne contributed to the taxonomy development and to the Ego-Exo Transfer baseline implementation and analysis. Avijit Dasgupta contributed to the annotation and taxonomy development, and to the early stage of baseline design. Miguel Martin contributed to the annotation and the taxonomy development. Shraman Pramanick contributed the EgoVLPv2 pretrained backbone. Yifei Huang contributed to the early stages of task definition and baseline design. Devansh Kukreja built the render flow to generate frame-aligned videos of each camera for each take as model input, and to produce video collages for annotations. Lorenzo Torresani contributed to editing this section. Kristen Grauman contributed to the task formulation.
13.B.2 Energy-efficient multimodal keystep recognition
Annotations and data
Formal task definition
In this task, the goal is to perform online classification of keysteps in a streaming egocentric multi-modal video, within an energy budget. We consider an ego video $\mathcal{V}_{ego}$ of arbitrary length $T$ comprising a stream of $K$ different sensory modalities (e.g., RGB images, audio, IMU, etc.). At each time step $t$, where $1\leq t\leq T$, the video consists of samples for each available modality, such that ${\mathcal{V}_{ego}^{t}=\{S^{t}_{1},\ldots,S^{t}_{K}\}}$, where $S^{t}_{j}$ denotes the sample at time $t$ for the $j^{\text{th}}$ modality. Given $\mathcal{V}_{ego}$ and an energy budget $B$, our task is to learn a model $\mathcal{F}$ that maximizes the overall keystep recognition performance across the full video while also ensuring that the combined energy for sensing and running model inference does not exceed $B$. $F$ consists of a sensor triggering policy $\mathcal{F}^{\mathcal{P}}$ and a keystep prediction model $\mathcal{F}^{\mathcal{K}}$. At every step $t$, the policy $\mathcal{F}^{\mathcal{P}}$ decides which sensor(s) to activate and sample from, in order to produce the model’s current observation ${O^{t}}$, such that $O^{t}\subseteq\{S^{t}_{1},\ldots,S^{t}_{K}\}$. Given $O^{t}$, the keystep predictor $\mathcal{F}^{K}$ outputs its estimate of the ground truth keystep for the current step.
The energy budget accounts for the cost of operations in each model forward pass, the cost of moving intermediate activations in and out of memory and the cost of the continuous operation of sensors, each having different cost profiles (e.g., IMU and audio sensors are relatively cheaper to operate than camera sensors). Note that the sensor triggering policy may be static (e.g., sample video at 4 frames per second (fps), keep audio/IMU off; sample 1 fps video, keep IMU always on) or dynamic (e.g., depending on the audio, decide whether to trigger video capture). We keep our task definition general, allowing the challenge to admit a wide variety of recent approaches ranging from pure video-based efficient backbone architectures [41] to multi-modal triggering approaches and, naturally, a combination of them.
Note that at test time, the input to the model can only include current and past observations as our task is strictly an online recognition task. However, we encourage exploring other modalities than those considered in our experiments, e.g., IMU or camera poses inferred from video, audio, and IMU.
Measuring energy consumption
Accurately measuring energy consumption of models is crucial for their use in AR/VR devices [2, 23]. The energy used comes from a complex interplay of sources including sensors, compute, communication, data processing, memory transfer (SRAM and DRAM), and leakage – many of which are typically ignored when building efficient computer vision models, despite their large energy consumption (e.g., memory transfer accounts for over 50% of the total power [183]).
We consider three key factors when modeling energy consumption following prior work [153]. (1) Compute energy: the cost of each model forward pass as a function of the number of operations (MACs). (2) Memory transfer energy: the cost associated with memory read-write operations for storing intermediate activations and model outputs. (3) Sensor triggering energy: the cost associated with turning on / off and continuous operation of sensors (camera, audio, IMU). For a model that processes an observation $O^{t}$, the total energy consumed can then be formulated as.
$E(O^{t})=\alpha*C(O^{t})+\beta*M(O^{t})+\sum_{j=1\ldots K}\gamma_{j}*\mathbbm{1}(S_{j}\in O^{t}),$ (1)
where $C(O^{t})$ corresponds to the total number of multiply-add operations computed during the forward pass (in MAC/s), $M(O^{t})$ corresponds to the total memory transferred to/from DRAM (in MB/s), and $S_{j}\in O^{t}$ corresponds to whether the $j$-th sensor is active. Finally, $\alpha,\beta,\gamma_{j}$ are weighting factors that measure the contribution of each energy source. We select these weighting parameters to reflect real-world AR/VR hardware capabilities. Namely, $\alpha$ = 4.6 pJ/MAC [153, 31]; $\beta$ = 80 pJ/byte [55]; $\gamma_{rgb}$ = 15 mW and $\gamma_{audio}$ = 0.5 mW [91].
While physical measurements are required for truly accurate energy estimates, this formula provides a reasonable approximation to it and can serve as a target for the computer vision community towards energy-efficient models for real-world devices.
@@ 13.B.2 Energy-efficient multimodal keystep recognition
Metrics
Following prior work [30], we evaluate online keystep detection performance using per-frame calibrated mean average precision (mcAP), which accounts for the imbalance in the keystep labels in our dataset. We measure energy consumption in mW as described above.
There is a natural trade-off between efficiency and better performance. Thus, we evaluate models in two tiers by setting a budget for the power consumption in each tier, namely 20 mW for the high-efficiency tier and 2.8W for the high-performance tier. We select the high-efficiency budget based on the energy consumption of current single-modality, efficient architectures (e.g., X3D-XS [41]) with an eye to the future where multi-modal models operate within it. For the high-performance tier, we set the budget to a value that permits the use of powerful transformer-based action recognition models like LaViLa [199]. Once a model runs out of budget, in our setup it uses its latest prediction for all future steps.
@@
We instantiate the task by considering keystep prediction episodes where the multimodal samples arrive in a streaming fashion. As mentioned above, we use vision and audio as our task modalities, where vision comprises RGB frames that are streaming at 30 frames per second (fps), and the audio modality is made up of time-aligned single-channel chunks that are 0.4 seconds long and sampled at 16 kHz. However, the setup can be extended to include IMU, and potentially other sensors as well. We evaluate all models at the rate of 5 fps on a total of 211 test episodes of variable length, where the shortest episode is $\sim$15 seconds, and the longest episode is $\sim$34 minutes. We filter out episodes where all steps belong to the background class and apply the same filtering when testing on v2. This leads to our test set being the same in both v1 and v2.
(a) High-efficiency tier (budget = 20 mW)
(b) High-performance tier (budget = 2825.71 mW)
Baselines
We provide a family of (less/more expensive) keystep prediction models for solving the task. Each model has a unimodal or audio-visual feature encoder followed by a keystep classification head.
  • X3D-XS [41]. This is a vision-only model comprising the X3D-XS feature encoder, which progressively expands the feature size and representational capacity of its layers, and later contracts them for achieving better performance-efficiency trade-off. This is the most lightweight model in our family of keystep predictors. The encoder has a depth factor of 2.2, and takes 4 RGB frames of size $160\times 160$ sampled at 15 fps, as inputs.
  • LaViLa [199]. This is another vision-only model where the visual feature encoder is trained through CLIP-style video-language pre-training. To improve the feature quality over vanilla CLIP-style pre-training, this method augments the number of video-text pairs by leveraging pre-trained large language models (LLM) to generate textual descriptions of un-annotated videos and rephrase existing narrations. In particular, we use the frozen TimeSformer [13]-Base (TSF-B) visual encoder pre-trained on the Ego4D dataset. To generate the feature for a target frame, the encoder samples 12 RGB frames of size $224\times 224$ at 30 fps from a time window centered around the target frame and pads the samples with the boundary frames on both ends to create a 16-frame clip.
  • Light-ASDNet [83]. This is an audio-only model that represents audio as spectrograms and efficiently encodes them by splitting 2D convolutions into 1D convolutions along the spectrogram temporal dimension [83]. In our setup, the spectrograms are Kaldi [122]-compliant, and consist of 196 temporal windows and 160 Mel-frequency bins, respectively.
  • Fixed stride. This is a policy that samples the input (video or audio) every $s$ prediction steps. We evaluate different $s$ values, where $s\in\{2,5,10,15,21,43,64,86,107,129,150\}$.
  • Audio-Visual (AV) Cascade. This is a policy that initially uses the Light-ASDNet model to predict the keystep, and switches over to the LaViLa model (expecting a more accurate prediction from it) if the audio-based prediction confidence is below a certain threshold. We set the confidence threshold to 0.5 in our experiments.
@@
We train all keystep prediction models using the data from v1 for 150 epochs using the cross entropy loss. We use the AdamW [95] optimizer with an initial learning rate of $10^{-4}$ and a weight decay of $10^{-5}$. We set the batch size to 512 for vision-only models, and 384 for audio-only and audio-visual models.
@@
Figure 39: Improvement (left) or degradation (right) in keystep recognition performance per keystep label, when comparing the most efficient vision-only (X3D-XS [41] + $s=10$) and audio-only (Light-ASDNet [83] + $s=5$) models from the high-efficiency tier. The plots show the 15 keysteps where improvement or degradation are largest. $\Delta$ reports the amount of improvement/degradation.
Method Modality mcAP (%) $\uparrow$ Power (mW) $\downarrow$
Light-ASDNet [83] + $s=5$ A 65.18 19.67
X3D-XS [41] + $s=10$ V 76.85 19.14
AV-LF w/ X3D-XS + $s=15$ AV 77.89 19.70
Table 15: Keystep prediction results for the high-efficiency tier (budget = 20 mW).
Method Modality mcAP (%) $\uparrow$ Power (mW) $\downarrow$
Lavila [199] + $s=5$ V 93.24 2245.66
AV-LF w/ Lavila + $s=5$ AV 92.18 2274.40
Table 16: Keystep prediction results for the high-performance tier (budget = 2.8W).
Results
In Fig. 38(a), we plot the recognition mcAP of all models against their total power consumption for the high-efficiency tier. We can see that combining vision and audio is better than using only vision or audio. Thus suggests that the two modalities carry complementary cues that are useful for the task. However, all vision-only models outperform their audio-only counterparts, which indicates that vision is the most critical modality for the task. The raw backbones generally perform better than the models using a sampling policy, but at the cost of requiring higher energy, making them impractical to use in online settings. Among the models that use a fixed stride, a lower stride generally improves the performance while hurting energy efficiency. For the vision-only and audio-only backbones, the best stride values within budget are $s=10$ and $s=5$, respectively. Among the audio-visual models, AV-LF with a stride of $s=15$ performs the best among all models that satisfy the budget. Using the greedy or random policy with AV-LF leads to a sharp decline in performance compared to using a fixed stride, showing that sampling very early or randomly in the episode is suboptimal for our online recognition task. AV-cascade also performs worse than most audio-visual models while also requiring more energy, possibly because the audio backbone often outputs wrong but over-confident predictions that prevent switching over to the more reliable vision backbone when required.
For easy reference, in Table 15 we report the recognition performance and total power consumption of our best uni-modal and audio-visual models within budget for the high-efficiency tier.
In Fig. 38(b), we plot the recognition mcAP of all models against their total power consumption for the high-performance tier. Different from the high-efficiency tier, the audio-visual backbone generally performs worse than the vision-only backbone, possibly because the LaViLa features are strong enough by themselves, and fusing them with audio features through the simple mechanism of linear late fusion reduces their expressivity. Otherwise, the overall behavior of different sampling policies is similar across the two tiers. We report the recognition performance and total power consumption of the best uni-modal and audio-visual models within the high-performance budget in table 16.
@@ Results
Additionally, in Fig. 39, we present a detailed analysis of the keystep labels where the best vision-only model from the high-efficiency tier yields the maximum improvement or decline in performance compared to its audio-only counterpart. We observe that the vision-only model produces a large improvement over the audio-only model usually in steps where the activity does not produce distinctive sounds (e.g., add green chillies, get celeries, etc.). On the other hand, using audio alone helps the most when the activities involve sounds that are strongly indicative of the nature of the task (e.g., stir fry egg mixture, cut butter, etc.).
@@
Finally, we envision that future work on this task will explore more sophisticated learned policies, potentially trained using reinforcement learning, in order to adaptively decide when to sample which modality instead of using fixed heuristics. Another promising direction is to investigate efficient transformer-based recognition backbones [180, 198] that can improve recognition performance without significantly affecting the model efficiency.
Contribution statement
Tushar Nagarajan led the energy-efficient multimodal benchmark, co-developed the task formulation, and advised the baseline development. Sagnik Majumder led the baseline development effort and contributed all baseline implementations and analysis of results for the benchmark. Merey Ramazanova developed the energy profiler used to evaluate all baselines and contributed to the experimental analysis. Lorenzo Torresani contributed to editing this section. Mitesh Kumar Singh helped design the energy formula. Miao Liu and Shengxin Cindy Zha initiated this benchmark and developed an early version of the task formulation.
13.B.3 Procedure understanding
The real-world motivation for our procedure understanding task has basis in both augmented reality (AR) and robotics. AR assistants, beyond recognizing the current keystep, could verify missing mandatory keysteps, suggest possible future ones, and detect procedural mistakes. Similarly, robots could learn the structure of a procedure from human demonstrations. Mining the structure of procedures has been shown useful for planning [21, 14] and improving keystep recognition [9, 203] and discovery [10].
Annotations
For each of the considered procedural tasks, we manually labeled task-graphs as structures encoding the keystep orderings leading to a correct execution of the procedure.
Figure 40: Example task-graph of a "Cooking Omelet" procedure.
Task-graphs. We define a task-graph as a directed graph in which nodes represent keysteps and directed edges represent dependencies. For instance, in the task-graph reported in Figure 40, the “Add Milk $\to$ Get a Bowl” structure denotes that keystep “Get a Bowl” has to be executed before keystep “Add Milk”. Besides directed edges, task-graphs also contain “OR” and “XOR” structures, which combine dependencies logically, as well as “optional” and “repeatable” node attributes.
Task-graph construction. We first familiarized ourselves with the procedural tasks by watching videos with annotated keysteps. We then initialized task graphs with procedural dependencies obtained from keystep annotations through the following procedure: a) a directed graph is first generated from the observed keystep transition frequencies; b) edges of the transition graph are filtered based on transition probabilities using a threshold parameter which is manually tuned for each scenario; c) edge directions are inverted to convert frequent transitions into dependencies. These initial graphs were then refined and corrected by human annotators leveraging common sense and in-domain knowledge.
Segment-level annotations. Let $S=\{s_{1},\ldots,s_{n}\}$ be a labeled sequence of keysteps in a given video. We denote with $y_{i}$ the annotated keystep label of segment $s_{i}$ and with $Y_{:i}=\{y_{1},\ldots,y_{i}\}$ the sequence of labels up to the $i$-th keystep. Using these keystep annotations, each segment $s_{i}$ is automatically matched to a task-graph and augmented with the following attributes: 1) a list of previous keysteps—these are the in-neighbors of the matched node, 2) optional labels—directly derived from the optional node attribute, 3) a procedural mistake label—this is set to “true” if the in-neighbors of the matched node do not correspond to segments in the history $Y_{:i}$, 4) the list of missing keysteps—the in-neighbors of the matched node not listed in $Y_{:i}$, and 5) the list of next steps—nodes for which in-neighbors appear in $Y_{:i}$. Non-repeatable nodes are listed only if they do not appear in $Y_{:i}$.
(a)Refer to caption (b)Refer to caption
Figure 41: Overview of the two procedure understanding approaches considered in our evaluation: (a) graph-based baselines for procedure understanding rely on a Keystep Assignment and a Procedural Reasoning component; (b) the architecture of our end-to-end baseline.
Formal task definition
Given a video segment $s_{i}$ and its segment history $S_{:i-1}=\{s_{1},\ldots,s_{i-1}\}$, models have to 1) determine previous keysteps (to be performed before $s_{i}$); infer if $s_{i}$ is 2) optional or 3) a procedural mistake; 4) predict missing keysteps (which should have been performed before $s_{i}$); and 5) forecast next keysteps (for which dependencies are satisfied and hence can be executed next). The task is weakly supervised, with two versions based on the level of supervision: 1) instance-level: segments and their keystep labels are available during training and inference; 2) procedure-level: unlabeled segments and procedure-specific keystep names are given for training and inference.
Note that, when the procedure-level supervision is considered, the input to the model excludes keystep labels both at training and test time. At both the procedure and instance levels of supervision, models are required to process the video in a causal fashion, meaning that predictions made at time $t$ only depend on observations made at time $t\textquoteright<t$.
Baselines
We consider two main baseline approaches: graph-based baselines, which leverage an explicit procedure structure mined from videos, and an end-to-end model aimed to make predictions directly from the observed videos.

### v4

@@ E.1.2 Ego-exo translation
In Table 20 we provide a break down of ego-exo translation results across different scenarios, using GNT-mask for track prediction and DiT-pix for clip generation. We can observe similar trends for the two subtasks: the methods achieve better results in basketball and soccer scenarios than in bike and cook scenarios, which is reasonable as the objects in bike and cook scenarios are more complex and diverse.
@@
Basketball 14.5 0.41
Soccer 17.0 0.51
Music 4.5 0.30
Health 12.9 0.40
Bike 6.4 0.52
Cook 9.5 0.47
Table 20: Breakdown of ego-exo translation results per scenario for the subtasks of ego track prediction (IoU) and ego clip generation (LPIPS).
E.2 Fine-grained keystep recognition
Data annotation details
@@ E.2 Fine-grained keystep recognition
We collect manual annotations for keysteps (actions that contribute towards the completion of a procedural task) and build a keystep taxonomy in parallel. Figure 40 shows the annotation user interface. We provide annotators a composite view of time-synchronized ego and exo videos. Each keystep annotation contains the start and end timestamps, a category label, a natural language description, and a flag indicating whether the keystep is essential or optional for task completion. Annotators interact with a search widget which displays keystep labels with their complete path within a hierarchical tree, e.g., Making cucumber & tomato salad $>$ Prepare dressing $>$ To a bowl or jar $>$ Add salt.
@@
Dataset splits
The keysteps in our dataset exhibit a very long-tailed distribution. To address this challenge, we set a cutoff threshold at 20 samples per keystep, limiting our analysis to 278 unique keysteps. For simplicity, we consider only the leaf node keysteps in the hierarchy. Exploring the hierarchical structure including parent nodes is a promising direction but we leave this as future work. In all, the dataset for keystep recognition comprises 130,979 segments, with an average duration of 11.34 seconds each. Specifically, the training set contains 74,342 segments, of which 14,326 are from the ego view and the rest from the exo view. The validation set consists of 23,636 segments, including 4,517 ego-view segments, and the test set has 33,001 segments with 6,373 in the ego view.
@@
Results breakdown by keystep
@@
We present a more detailed analysis of per-step performance in Figure 42, comparing training with ego-view videos and exo-view videos. We can observe that exo views show performance advantages over ego views in several steps, with the keystep ‘have a conversation asking different questions’ benefiting the most from exo. Conversely, ego views are more effective in steps involving manipulation of small objects, like ‘cut carrots’ and ‘unpack the new tube’. This observation can be linked to the positioning of exo cameras, which are often placed further away from the subject, enabling them to capture a broader view, though possibly missing finer details of the activity. We hope these findings provide insight for future research on the effective use of exo-view videos during training.
@@
Figure 42: Keystep recognition evaluated per keystep label on a validation split, comparing training with only ego-view videos versus exo-view videos. The accuracy delta (exo-ego) is displayed, where a positive value indicates better exo and a negative value indicates better ego performance.
E.3 Energy-efficient multimodal keystep recognition
@@
Energy tiers
As mentioned in the main paper, there is a natural trade-off between efficiency and better performance. Thus, we evaluate models in two tiers by setting a budget for the power consumption in each tier, namely 20 mW for the high-efficiency tier and 2.8W for the high-performance tier. We select the high-efficiency budget based on the energy consumption of current single-modality, efficient architectures (e.g., X3D-XS (Feichtenhofer, 2020)) with an eye to the future where multi-modal models operate within it. For the high-performance tier, we set the budget to a value that permits the use of powerful transformer-based action recognition models like LaViLa (Zhao et al., 2023). Once a model runs out of budget, in our setup it uses its latest prediction for all future steps.
Complete baseline details
  • X3D-XS (Feichtenhofer, 2020). This is a vision-only model comprising the X3D-XS feature encoder, which progressively expands the feature size and representational capacity of its layers, and later contracts them for achieving better performance-efficiency trade-off. This is the most lightweight model in our family of keystep predictors. The encoder has a depth factor of 2.2, and takes 4 RGB frames of size $160\times 160$ sampled at 15 fps, as inputs.
  • LaViLa (Zhao et al., 2023). This is another vision-only model where the visual feature encoder is trained through CLIP-style video-language pre-training. To improve the feature quality over vanilla CLIP-style pre-training, this method augments the number of video-text pairs by leveraging pre-trained large language models (LLM) to generate textual descriptions of un-annotated videos and rephrase existing narrations. In particular, we use the frozen TimeSformer (Bertasius et al., 2021)-Base (TSF-B) visual encoder pre-trained on the Ego4D dataset. To generate the feature for a target frame, the encoder samples 12 RGB frames of size $224\times 224$ at 30 fps from a time window centered around the target frame and pads the samples with the boundary frames on both ends to create a 16-frame clip.
  • Light-ASDNet (Liao et al., 2023). This is an audio-only model that represents audio as spectrograms and efficiently encodes them by splitting 2D convolutions into 1D convolutions along the spectrogram temporal dimension (Liao et al., 2023). In our setup, the spectrograms are Kaldi (Povey et al., 2011)-compliant, and consist of 196 temporal windows and 160 Mel-frequency bins, respectively.
@@
We instantiate the task by considering keystep prediction episodes where the multimodal samples arrive in a streaming fashion. As mentioned above, we use vision and audio as our task modalities, where vision comprises RGB frames that are streaming at 30 frames per second (fps), and the audio modality is made up of time-aligned single-channel chunks that are 0.4 seconds long and sampled at 16 kHz. However, the setup can be extended to include IMU, and potentially other sensors as well. We evaluate all models at the rate of 5 fps on a total of 211 test episodes of variable length, where the shortest episode is $\sim$15 seconds, and the longest episode is $\sim$34 minutes. We filter out episodes where all steps belong to the background class.
@@
We train all keystep prediction models for 150 epochs using the cross entropy loss. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with an initial learning rate of $10^{-4}$ and a weight decay of $10^{-5}$. We set the batch size to 512 for vision-only models, and 384 for audio-only and audio-visual models.
@@
Figure 43: Improvement (left) or degradation (right) in keystep recognition performance per keystep label, when comparing the most efficient vision-only (X3D-XS (Feichtenhofer, 2020) + $s=10$) and audio-only (Light-ASDNet (Liao et al., 2023) + $s=5$) models from the high-efficiency tier. The plots show the 15 keysteps where improvement or degradation are largest. $\Delta$ reports the amount of improvement/degradation.
Performance breakdown by modality
@@
In Fig. 43, we present a detailed analysis of the keystep labels where the best vision-only model from the high-efficiency tier yields the maximum improvement or decline in performance compared to its audio-only counterpart. We observe that the vision-only model produces a large improvement over the audio-only model usually in steps where the activity does not produce distinctive sounds (e.g., add green chillies, get celeries, etc). On the other hand, using audio alone helps the most when the activities involve sounds that are strongly indicative of the nature of the task (e.g., stir fry egg mixture, cut butter, etc).
@@
Finally, we envision that future work on this task will explore more sophisticated learned policies, potentially trained using reinforcement learning, in order to adaptively decide when to sample which modality instead of using fixed heuristics. Another promising direction is to investigate efficient transformer-based recognition backbones (Xu et al., 2021, Zhao and Krähenbühl, 2022) that can improve recognition performance without significantly affecting the model efficiency.
E.4 Procedure Understanding
Complete baseline details